# 1st Pranav Kulkarni    2nd Leo Sünkel    3rd Michael Kölle

### v2

@@ IV-D Continuous Variable Distillation
For continuous variable entanglement, APC implements an effective model for applying an NLA-based distillation stage to a two-mode squeezed state subjected to symmetric loss. The backend tracks an effective squeezing parameter $r$ (or equivalently $\lambda=\tanh r$) and a loss parameter $\eta$, and maps these to a fidelity proxy $F_{\text{CV}}$ with respect to an ideal entangled Gaussian resource via a monotonically increasing function of $r$ and $\eta$ [16, 17]. A sequence of $K$ NLA stages with gain $g$ yields an updated effective squeezing $r^{\prime}_{\text{eff}}$ and a success probability of the form
@@ IV-E1 Per-Link Design Space and Pareto Frontiers
Figure 2: APC rounds adapt to gate noise on a 3-hop chain. Setup: total 24 km (3 links), $F_{0}=0.93$ per link giving $F_{\mathrm{raw}}\approx 0.80$ after two swaps, target $F^{\star}=0.85$, gate or measurement error probability $\varepsilon$ swept from $10^{-4}$ to $3\times 10^{-2}$. (a) End-to-end fidelity versus $\varepsilon$ (semi-log in $\varepsilon$) with threshold $F^{\star}$ marked. (b) Goodput versus $\varepsilon$ (log-log); APC sustains $\sim 10^{2}$ to $5\times 10^{2}$/s until the noise cliff. (c) APC-selected rounds (1 to 4) versus $\varepsilon$, compensating for increasing noise until saturation. (d) At high noise beyond the cliff, additional rounds reduce fidelity because CNOT-induced errors dominate; $r=0$ becomes least bad.
@@
Figure 8: APC planning is fast enough for real-time control. (a) Planning time versus chain length for 1 to 1000 links (three calls averaged per point) with $F_{0}=0.88$, $g_{\mathrm{2q}}=0.9999$, $F^{\star}=0.85$, $T_{2}=1.0$ s; growth is roughly linear and remains near 6–7 s at 1000 links on an M1 Pro (Apple Silicon). (b) Total CPU time to plan $m$ independent single-link requests for $m\in\{1,4,8,16,32,64\}$; sequential planning scales linearly, indicating parallelization headroom. (c) CPU planning time versus theoretical protocol makespan (log scale) across chain length; planning is negligible for small chains and remains small compared to execution even at large scale. (d) Amortized planning time per link decreases with chain length, approaching a few milliseconds per link around 1000 links.
@@ IV-E3 Objectives and Goodput
Because static purification schemes that apply the same protocol and depth on every hop correspond to restricting $\mathcal{P}_{\ell_{1}}=\cdots=\mathcal{P}_{\ell_{H}}=\{u_{\text{static}}\}$, they are represented as a subset of APC’s design space. For any additive or monotone objective $J$, the global optimum over APC’s full design space is therefore never worse than the optimum over static schemes; this is formalized in Section VI.
@@ VI Results and Discussion
VI-A Protocol Identity versus Depth (Fig. 5)
Under Werner noise, BBPSSW and DEJMPS provide nearly identical fidelity gains per round. The APC’s advantage therefore comes from choosing the right number of rounds, not the protocol identity. As $F^{\star}$ increases, the APC climbs in $r$ while single-protocol static strategies either waste effort by overshooting or fail the constraint.
VI-B Maximizing Goodput via Adaptive Depth
@@
VI-C Feasibility Envelope and Noise Robustness
The operational limits of the network are defined by the “feasibility envelope”, the region in parameter space where $F_{\text{end}}\geq F^{\star}$ is physically achievable. Fig. 3 delineates this boundary across link distance and target fidelity. We observe a sharp cutoff where the combination of transmission loss (lowering rate) and classical signaling latency (increasing decoherence) renders high-fidelity distribution impossible, regardless of the purification depth. This limitation is further dissected in Fig. 2, which isolates the impact of gate errors. While purification can suppress channel noise, it accumulates local gate errors. Consequently, we observe a “noise cliff” at gate error rates $\approx 10^{-3}$. Beyond this threshold, aggressive purification ($r\geq 2$) becomes detrimental, as the entropy added by noisy CNOT gates outweighs the distillation gain. The APC correctly identifies this regime (Fig. 2c), truncating the purification depth to maximize utility in high-noise environments.
VI-D Impact of Memory Coherence
A critical constraint in wide-area quantum networks is memory coherence time ($T_{2}$). As shown in Fig. 4, there exists a minimum $T_{2}$ required to support any purification at all. For a 3-hop chain targeting $F^{\star}=0.76$, coherence times below $10$ ms render the protocol infeasible (Fig. 4a). This is because the classical round-trip time required to herald purification outcomes allows stored qubits to decohere below the distillation threshold. The APC adapts to this by disabling purification ($r=0$) in low-$T_{2}$ regimes (Fig. 4b), effectively reverting to a “repeat-until-success” swapping strategy to salvage connectivity, albeit at lower fidelity.
VI-E Dynamic Protocol Selection
@@
VI-F Multipartite and Continuous Variable Extensions
We extended the evaluation to multipartite GHZ state distribution (Fig. 6) and continuous variable entanglement (Fig. 7). For GHZ states, the goodput scales as $\sim p_{\text{link}}^{k}$, resulting in an exponential decay with the number of parties $k$ (Fig. 6a). The APC mitigates this by optimizing the underlying bipartite links, but the results highlight the necessity of high-efficiency fusion gates for scaling beyond $k=5$. For CV systems, we simulated Noiseless Linear Amplification (NLA). Fig. 7 confirms that optimal gain choices ($g\approx 1.1-1.5$) can counteract channel loss, though the probabilistic nature of NLA imposes a heavy rate penalty similar to discrete variable recurrence.
VI-G Real-Time Planning Latency
Finally, we benchmarked the computational overhead of the APC planner itself (Fig. 8). The dynamic programming algorithm exhibits near-linear scaling with chain length. For a 1000-link chain, the planning latency is approximately 6–7 seconds on a M1 Pro processor, with a per-link amortized cost of milliseconds. This confirms that the APC is computationally lightweight enough to operate in the control plane of a software-defined quantum network, calculating optimal policies on timescales orders of magnitude faster than the drift of physical hardware parameters.

### v3

@@ IV-D Continuous Variable Distillation
For continuous variable entanglement, APC implements an effective model for applying an NLA-based distillation stage to a two-mode squeezed state subjected to symmetric loss. The backend tracks an effective squeezing parameter $r$ (or equivalently $\lambda=\tanh r$) and a loss parameter $\eta$, and maps these to a fidelity proxy $F_{\text{CV}}$ with respect to an ideal entangled Gaussian resource via a monotonically increasing function of $r$ and $\eta$ [11]. A sequence of $K$ NLA stages with gain $g$ yields an updated effective squeezing $r^{\prime}_{\text{eff}}$ and a success probability of the form
@@ IV-E1 Per-Link Design Space and Pareto Frontiers
Figure 2: APC rounds adapt to gate noise on a 3-hop chain. Setup: total 24 km (3 links), $F_{0}=0.93$ per link giving $F_{\mathrm{raw}}\approx 0.80$ after two swaps, target $F^{\star}=0.85$, gate or measurement error probability $\varepsilon$ swept from $10^{-4}$ to $3\times 10^{-2}$. (a) End-to-end fidelity versus $\varepsilon$ with threshold $F^{\star}$ marked (to reduce parameter space, we use a correlated noise sweep $p_{1}=p_{2}=p_{\mathrm{meas}}=\varepsilon$). (b) Goodput versus $\varepsilon$ (log-log); APC sustains $\sim 10^{2}$ to $5\times 10^{2}$/s until the noise cliff. (c) APC-selected rounds (1 to 4) versus $\varepsilon$, compensating for increasing noise until saturation. (d) At high noise beyond the cliff, additional rounds reduce fidelity because CNOT-induced errors dominate; $r=0$ becomes least bad.
@@
Figure 8: APC planning is fast enough for real-time control. (a) Planning time versus chain length for 1 to 1000 links (three calls averaged per point) with $F_{0}=0.88$, $p_{q}=0.9999$, $F^{\star}=0.85$, $T_{2}=1.0$ s; growth is roughly linear and remains near 6–7 s at 1000 links on an M1 Pro (Apple Silicon). (b) Total CPU time to plan $m$ independent single-link requests for $m\in\{1,4,8,16,32,64\}$; sequential planning scales linearly, indicating parallelization headroom. (c) CPU planning time versus theoretical protocol makespan (log scale) across chain length; planning is negligible for small chains and remains small compared to execution even at large scale. (d) Amortized planning time per link decreases with chain length, approaching a few milliseconds per link around 1000 links.
@@ IV-E3 Objectives and Goodput
Because static purification schemes that apply the same protocol and depth on every hop correspond to restricting $\mathcal{P}_{\ell_{1}}=\cdots=\mathcal{P}_{\ell_{H}}=\{u_{\text{static}}\}$, they are represented as a subset of APC’s design space. For any additive or monotone objective $J$, the global optimum over APC’s full design space is therefore never worse than the optimum over static schemes; this is illustrated empirically in Section VI.
@@ VI Results and Discussion
VI-A Maximizing Goodput via Adaptive Depth
@@
VI-B Feasibility Envelope and Noise Robustness
The operational limits of the network are defined by the “feasibility envelope”, the region in parameter space where $F_{\text{end}}\geq F^{\star}$ is physically achievable. Fig. 3 delineates this boundary across link distance and target fidelity. We observe a sharp cutoff where the combination of transmission loss (lowering rate) and classical signaling latency (increasing decoherence) renders high-fidelity distribution impossible, regardless of the purification depth. This limitation is further dissected in Fig. 2, which isolates the impact of gate errors. While purification can suppress channel noise, it accumulates local gate errors. Here, we use a correlated noise sweep $p_{1}=p_{2}=p_{\mathrm{meas}}=\varepsilon$ to reduce parameter space (results should be interpreted as joint scaling of local error sources). Consequently, we observe a “noise cliff” at gate error rates $\approx 10^{-3}$. Beyond this threshold, aggressive purification ($r\geq 2$) becomes detrimental, as the entropy added by noisy CNOT gates outweighs the distillation gain. The APC correctly identifies this regime (Fig. 2c), truncating the purification depth to maximize utility in high-noise environments.
VI-C Impact of Memory Coherence
A critical constraint in wide-area quantum networks is memory coherence time ($T_{2}$). As shown in Fig. 4, there exists a minimum $T_{2}$ required to meet the fidelity constraint under the end-to-end timing model (and, if needed, enable recurrence). For a 3-hop chain targeting $F^{\star}=0.76$, coherence times below $10$ ms render the protocol infeasible (Fig. 4a). This is because the classical round-trip time required to herald purification outcomes allows stored qubits to decohere below the distillation threshold. The APC adapts to this by disabling purification ($r=0$) in low-$T_{2}$ regimes (Fig. 4b), effectively reverting to a “repeat-until-success” swapping strategy to salvage connectivity, albeit at lower fidelity.
VI-D When protocol identity matters
Fig. 5–Under Werner noise, BBPSSW and DEJMPS provide nearly identical fidelity gains per round. The APC’s advantage therefore comes from choosing the right number of rounds, not the protocol identity. As $F^{\star}$ increases, the APC climbs in $r$ while single-protocol static strategies either waste effort by overshooting or fail the constraint.
@@ VI-D When protocol identity matters
VI-E Multipartite and Continuous Variable Extensions
We extended the evaluation to multipartite GHZ state distribution (Fig. 6) and continuous variable entanglement (Fig. 7). For GHZ states, the goodput scales as $\sim p_{\text{link}}^{k}$ (the per-link success probability after purification), resulting in an exponential decay with the number of parties $k$ (Fig. 6a). The APC mitigates this by optimizing the underlying bipartite links, but the results highlight the necessity of high-efficiency fusion gates for scaling beyond $k=5$. For CV systems, we simulated Noiseless Linear Amplification (NLA). Fig. 7 confirms that optimal gain choices ($g\approx 1.1-1.5$) can counteract channel loss, though the probabilistic nature of NLA imposes a heavy rate penalty similar to discrete variable recurrence.
VI-F Real-Time Planning Latency
Finally, we benchmarked the computational overhead of the APC planner itself (Fig. 8). The dynamic programming algorithm exhibits near-linear scaling with chain length. For a 1000-link chain, the planning latency is approximately 6–7 seconds on a M1 Pro processor, with a per-link amortized cost of milliseconds. This confirms that the APC is computationally lightweight enough to operate in the control plane of a software-defined quantum network, calculating optimal policies on timescales faster than the drift of physical hardware parameters.